R2 = 0.5754 and n = 425. The chi-square statistic for testing the overall regression effect H0 for both slopes is 0. 575.9421: How many degrees of freedom are there for the chi-square test statistic?
Answer #1

degree of freedom in chi square test = (no. of rows - 1)*(no. of columns - 1)

no. of rows = n = 425

no. of columns in regreesion = 2 (in regression there are 2 columns x and y)

degree of freedom = (425-1)*(2-1) = 424

degree of freedom = 424

(please upvote)